OpenStudy (anonymous):

A wooden plate is shaped as a pentagon. The measure of one angle is 40°. The remaining interior angles are obtuse and of equal measure. What is the measure of each obtuse angle in the pentagon?

OpenStudy (anonymous):

A. 120° B. 125° C. 110° D. 100°

OpenStudy (anonymous):

@UnkleRhaukus @robtobey Do You All Under Stand This Question?

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

Do you use the Interior Angle Sum Theorem? @UnkleRhaukus

OpenStudy (unklerhaukus):

that is a good idea

OpenStudy (anonymous):

But It Doesn't Give Me The Answer In The Answer Choices So PLEASE HELP ME! @UnkleRhaukus

OpenStudy (unklerhaukus):

what will be the interior angle sum of a pentageon?

OpenStudy (anonymous):

540? @UnkleRhaukus

OpenStudy (unklerhaukus):

right so you can set up an equation 540°= 40°+4x°

OpenStudy (anonymous):

So 540-40= 500/4=125 So The Answer Is 125 ^_^

OpenStudy (anonymous):

THANK YOU! @UnkleRhaukus